# Riesz Representer Fitting under Bregman Divergence: A Unified Framework for Debiased Machine Learning

### v1

@@ Appendix G Riesz Regression and Density Ratios
As explained in the main text and in Kato (2025b), the Riesz representer is closely connected to density ratio estimation. In particular, for ATE, the Riesz representer can be expressed in terms of two density ratios relative to the marginal covariate distribution, which leads to a decomposition of the squared loss objective into two LSIF problems.
@@ H.1 Nearest Neighbor Matching
Following this study, Kato (2025a) shows that nearest neighbor matching for ATE estimation can be interpreted as a special case of SQ-Riesz regression, that is, Riesz regression or LSIF. The key step is to express the ATE Riesz representer $\alpha^{\text{ATE}}_{0}(D,Z)$ in terms of density ratios with respect to the marginal covariate distribution, and to approximate these density ratios via nearest neighbor cells, following the density ratio interpretation in Lin et al. (2023).
@@ H.3 AME Estimation by Score Matching
A subsequent work Kato (2025c) shows that, for derivative-type linear functionals, the Riesz representer can be estimated via score matching. This principle also underlies score-based diffusion models (Song & Ermon, 2020; Song et al., 2021). This viewpoint is useful for AME and APE estimation, and for mitigating overfitting in flexible Riesz representer models, because score matching objectives introduce smoothing through derivatives or noise perturbations.
@@ H.4 Riesz Representer Estimation via Infinitesimal Classification
Next, following Kato (2025c), we introduce Riesz representer estimation via infinitesimal classification, which also reduces to score matching. This approach applies to a broader range of applications, not only to AME estimation.
@@
Kato (2025c) extends density ratio estimation via infinitesimal classification to Riesz representer estimation. In this subsection, we introduce an example of the method for APE estimation. For implementations in other applications, see Kato (2025c).

### v2

@@ Appendix G Riesz Regression and Density Ratios
As explained in the main text and in Kato (2025d), the Riesz representer is closely connected to density ratio estimation. In particular, for ATE, the Riesz representer can be expressed in terms of two density ratios relative to the marginal covariate distribution, which leads to a decomposition of the squared loss objective into two LSIF problems.
@@ H.1 Nearest Neighbor Matching
Following this study, Kato (2025c) shows that nearest neighbor matching for ATE estimation can be interpreted as a special case of SQ-Riesz regression, that is, Riesz regression or LSIF. The key step is to express the ATE Riesz representer $\alpha^{\text{ATE}}_{0}(D,Z)$ in terms of density ratios with respect to the marginal covariate distribution, and to approximate these density ratios via nearest neighbor cells, following the density ratio interpretation in Lin et al. (2023).
@@ H.3 AME Estimation by Score Matching
A subsequent work Kato (2025e) shows that, for derivative-type linear functionals, the Riesz representer can be estimated via score matching. This principle also underlies score-based diffusion models (Song & Ermon, 2020; Song et al., 2021). This viewpoint is useful for AME and APE estimation, and for mitigating overfitting in flexible Riesz representer models, because score matching objectives introduce smoothing through derivatives or noise perturbations.
@@ H.4 Riesz Representer Estimation via Infinitesimal Classification
Next, following Kato (2025e), we introduce Riesz representer estimation via infinitesimal classification, which also reduces to score matching. This approach applies to a broader range of applications, not only to AME estimation.
@@
Kato (2025e) extends density ratio estimation via infinitesimal classification to Riesz representer estimation. In this subsection, we introduce an example of the method for APE estimation. For implementations in other applications, see Kato (2025e).